# SkillRT: Compiling Skills for Efficient Execution Everywhere

### 03 Apr 2026 v1

@@ Abstract.
To address this challenge, we analyze  118,000 skills and draw inspiration from traditional compiler design. We treat skills as code and LLMs as heterogeneous processors. To make portability actionable, we decompose a skill’s requirements into a set of primitive capabilities, and measure how well each model-harness pair supports them. Based on these capability profiles, we propose SkillRT, a compilation and runtime system designed for portable and efficient skill execution. At compile time, SkillRT performs capability-based compilation, environment binding, and concurrency extraction. At runtime, SkillRT applies JIT code solidification and adaptive recompilation for performance optimization.
@@ 1. Introduction
Inspired by how classical computing systems handle code, we introduce SkillRT, a compilation and runtime system for skills that enables efficient cross-model execution and provides unified underlying support for skill execution. Through SkillRT, skills are compiled into forms tailored to different models, allowing each model to better understand and execute skills. Furthermore, SkillRT provides a unified runtime environment that systematically manages skill loading, parsing, and concurrent execution.
@@
We structure SkillRT around classical compilation techniques: interpreted execution (jansen2007interpretation, ), ahead-of-time (AOT) compilation (serrano2021javascript, ; serrano2018javascript, ), and just-in-time (JIT) optimization (aho2006compilers, ; aycock2003jit, ; suganuma2000overview, ; ishizaki2000study, ; cramer1997compiling, ). Currently, agents handle skills using only interpreted execution, feeding raw skill text directly to the model. This approach hampers skill portability and execution efficiency. In contrast, SkillRT further applies AOT and JIT compilation to optimize for diverse backend models and execution environments:
@@
AOT compilation: Skills inherently encode requirements of model capabilities, environment configuration, and parallelism opportunities. Therefore, SkillRT analyzes these requirements before execution and optimizes skills accordingly. Capability-based compilation characterizes the capability gap between LLMs and skill requirements by extracting 26 primitive capabilities. Each capability abstracts a distinct dimension of model behavior with multiple proficiency levels. The compiler measures the target model against these capability dimensions and adapts the skill specification to better align with model strengths and limitations. Second, environment binding extracts implicit dependencies on packages and tools from skill descriptions and generates setup scripts executable at load time, ensuring all dependencies are satisfied before skill execution. Finally, concurrency extraction draws inspiration from classical compiler optimizations for data-level parallelism (DLP) (kalathingal2016dynamic, ), instruction-level parallelism (ILP) (ranganathan1998empirical, ; kastner2001ilp, ), and thread-level parallelism (TLP) (redstone2003mini, ). It extracts parallelism opportunities at similar three granularities from skills and explicitly exposes them to the agent harness, thereby improving task execution efficiency.
@@
JIT optimization: At runtime, SkillRT leverages runtime information to continuously optimize skill performance and correctness. Code solidification: Skills may contain parameterized script templates which may be instantiated multiple times during different executions. JIT compilation materializes these high-frequency templates into instantiated executable code, bypassing LLM parsing. Adaptive recompilation: JIT monitoring tracks model execution and adaptively recompiles skills when capability gaps emerge, ensuring skills can self-improve through iterative optimization.
@@ 3. Design Overview
SkillRT is a compilation and runtime system for skills. It makes skills portable across heterogeneous targets by combining an AOT compiler (aho2006compilers, ) that specializes skills at install time with a runtime that resolves execution-time uncertainty. Figure 5 shows the architecture.
Figure 5. SkillRT architecture. The AOT compiler produces optimized skill variants at install time through three passes. The runtime manages variant selection, JIT optimization, and resource-aware scheduling during execution.
@@ 4. Skill Compilation
When a user installs a skill for the first time, SkillRT compiles it against the target (model, harness, host environment) before the skill ever runs. The compiler reads the raw skill text, identifies the target, and produces one or more optimized skill variants through three sequential passes. Each pass closes a different kind of gap between what the skill expects and what the target provides.
@@ 4.1. Capability-Based Compilation
The first pass addresses the model and harness mismatch. Because targets vary widely (§2.5), the compiler cannot embed optimization logic specific to each target. Instead, SkillRT introduces primitive capabilities, an abstract vocabulary for expressing what a skill requires and what a target can provide. The compiler profiles each target offline with microbenchmarks for these capabilities, compares the profiling result with the skill’s capability requirements, and selects the optimization strategy based on the gap.
@@ 4.1.2. Measuring the Gap Between Skill and Target
On the target side, SkillRT profiles the target with microbenchmarks generated for each primitive capability to measure how well the target supports it. Success on a microbenchmark for a primitive capability at a given level indicates that the model has attained at least that level of proficiency for that capability. This profiling is performed when the user sets up the agent and model. The results are cached and reused for all skill compilations for the same target.
@@ 4.3. Concurrency Extraction
The compiler uses LLM-assisted analysis to decompose the skill into discrete steps, identifying for each step what inputs it consumes and what outputs it produces. It then builds a workflow DAG: nodes are steps, and an edge from step A to step B exists when B consumes an output of A. The compiler also analyzes each step internally for sub-operations that can proceed concurrently. Figure 7 illustrates this process on an incident-response skill. For each parallel opportunity, whether between steps or within a step, the compiler selects a concurrency primitive based on the task structure and the harness. SkillRT defines three levels of parallelism.
@@ 5. Skill Runtime with JIT Optimization
AOT compilation generates multiple skill variants for different targets. Consequently, at runtime, SkillRT loads the skill variant compiled for the current target. Since AOT compilation is performed only once per skill and target (model+harness) pair, the compilation overhead is modest.
However, some problems only surface at execution time, such as capability gaps that static profiling missed or resource contention among parallel sub-agents competing for rate-limited APIs. Other opportunities, like code patterns that recur across invocations, only become visible after repeated runs. The SkillRT runtime addresses both through JIT optimization (§5.2) and a resource-aware scheduler (§5.3).
@@ 5.2. JIT Optimization
Beyond AOT compilation, which tailors primitive capabilities to each target, SkillRT further applies JIT optimization to enhance execution quality and efficiency through two complementary mechanisms: adaptive recompilation and code solidification.
@@ 5.2.1. Adaptive Recompilation
Figure 9. Skill compilation effectiveness. Cell values represent task completion rates after SkillRT optimization, while cell color indicates the score improvement (green) or regression (red) relative to original baseline skills.
@@ 5.2.2. Code Solidification
In the third stage, the template is instantiated into a callable function: the template code is wrapped with parameter extraction and output handling to produce a standalone shell script or code function. This instantiation step is lightweight and mechanical, as the AOT stage has already done the analysis. Subsequent invocations bypass the LLM entirely: the runtime extracts parameters from the task context, calls the instantiated function, and returns the result. If the solidified code execution causes an agent task to fail or encounters a runtime exception, SkillRT triggers a fallback mechanism that re-enables LLM-based code generation to ensure correctness.
@@ 5.3. Resource-Aware Parallel Scheduling
SkillRT dynamically monitors system state and schedules task execution accordingly, bridging the gap between compile-time opportunity and run-time reality. For API-bound workloads, it watches response latencies and HTTP 429 (rate limit) signals. For compute-bound workloads, it monitors CPU and memory usage. When pressure rises above a threshold, SkillRT applies two mechanisms to mitigate resource contention. First, it throttles the launch of new sub-agents to avoid introducing additional concurrency pressure. Second, for sub-agents that are already running, SkillRT selectively suspends a subset of them (i.e., pauses their agent-loop execution), based on per-agent resource usage or launch order, to reduce active competition for shared resources. SkillRT also records the effective concurrency of each skill from its previous run on the current system and uses it as a concurrency hint for subsequent executions.
@@ 6. Evaluation
Figure 10. Average task score by skill variant. Four variants are compared across eight models and three harnesses: No Skill, Original, Skill-Creator, and SkillRT-Optimized. Tier boundaries (SOTA / Mid / Small) are shaded. SkillRT-Optimized skills consistently achieve the highest scores, with the largest absolute gains for weaker models and on OpenClaw.
@@ 6.1. Setup
Models. We evaluate SkillRT across eight models spanning three capability tiers to comprehensively assess its effectiveness across varying model capacities. SOTA: claude-opus-4.6 (claude4_6_opus, ), deepseek-v3.2 (deepseek2025v3, ). Mid-tier: gemini-3-flash (gemini_flash, ), qwen3.5-397b (qwen3_5_397b, ), qwen3.5-122b (qwen3_5_122b, ), claude-3.5-haiku (claude3_5_haiku, ). Small: qwen3-30b (qwen3_30b, ), devstral-small (devstral_small_2507, ).
Baselines. We compare SkillRT against three baselines. No Skill: the agent executes without any skill loaded. Original: the original skill. Skill-Creator: the skill optimized by Anthropic’s Skill-Creator (anthropic-skill-creator, ) using claude-opus-4.6.
Harnesses. We select three different open-source agent harnesses to evaluate SkillRT performance across diverse execution environments. Each harness offers distinct functionality: BareAgent: a minimal harness that injects the compiled skill directly into the system prompt without additional abstraction layers. OpenCode (opencode, ): a full-featured code agent harness with batch tool dispatch capabilities and a rich tool set for code execution. OpenClaw (openclaw, ): a general-purpose agent harness with integrated functional modules designed to support complex multi-step task execution.
Methodology. We first profile the primitive capabilities of different models, and then apply SkillRT’s AOT and JIT compilation to generate optimized skill variants tailored to each model. To support the compiled skill artifacts, we extend the agent harnesses with custom loading and management mechanisms that enable SkillRT’s optimizations. For each task, we generate five diverse input instances and measure the average task completion rate, token consumption, and end-to-end execution latency. All experiments are conducted on a Mac Mini M4 with 16GB memory.
@@ 6.2. Skill Compilation Effectiveness
We first evaluate SkillRT’s effectiveness in improving task completion rates. We compare SkillRT-optimized skills against unoptimized baseline skills across different tasks, models, and agent harnesses. For SkillRT, the primary gains stem from capability-based compilation and JIT optimization. Figure 9 presents compilation effectiveness across eight models and three harnesses. Cell values represent task completion rates after SkillRT optimization, while cell color indicates the score improvement (green) or regression (red) relative to unoptimized baseline skills.
Figure 9 demonstrates SkillRT’s optimization effectiveness (compared with original skills) across diverse models and agent harnesses. Weaker models benefit more substantially from SkillRT optimization. This finding indicates that for most agent tasks, weaker models possess sufficient capability for the logical components but lack proficiency in non-logical aspects, such as generating complex JSON structures or managing environment dependencies. SkillRT’s compilation optimization addresses these gaps through primitive capability refinement, thereby improving overall task completion rates. For stronger models, improvements are more modest, as these models already handle tasks competently; SkillRT optimization primarily reduces token consumption and execution latency. Figure 10 aggregates average task scores across all four skill variants for each model and harness. SkillRT-optimized skills achieve the highest scores on every model–harness combination. The improvement over the Skill-Creator baseline grows with decreasing model capability: on BareAgent, SkillRT-optimized skills improve over Skill-Creator by 25% for qwen3-30b and 10% for devstral-small. In terms of regressions, SkillRT-compiled skills reduce task completion rates on only 4.5% of tasks, compared with 15% for the original skills. Using the original skill, OpenCode and OpenClaw differ by up to 13 points, depending on the model, which shows that harness behavior alone materially changes outcomes. SkillRT optimization raises scores on both harnesses and reduces this cross-harness gap to at most 5 points.
@@ 6.3. Staged Optimization Breakdown
In 11 of 14 categories, the original skill performs worse than using no skill. After SkillRT’s AOT compilation, the average task score improves by 88%. JIT optimization identifies additional capability defects that were not exposed during AOT compilation, further improving performance. After one JIT round, 8 of 14 skills achieve full scores on the evaluation tasks, and after three rounds, 10 of 14 do so.
@@ 6.4. Token and Cost Efficiency with Compilation
Beyond improving task success rates, SkillRT provides substantial cost benefits for SOTA LLMs by reducing token consumption. LLMs can correct execution errors through iterative agent loops, refining actions based on environment feedback. Therefore, even when tasks eventually succeed, this approach may incur substantial token waste. Figure 12 shows how task success rates and token costs vary across different model tiers and harnesses after SkillRT compilation.
Our evaluation reveals that for most models and harness combinations, SkillRT compilation simultaneously improves task success and reduces token consumption. For the strongest model and weakest harness pairing (DS-v3.2 + BareAgent), we observe token savings approaching 40%. This improvement arises because agent loops incur interaction overhead when correcting environment-related and tool-invocation errors; multiple retries are often required. By employing JIT compilation, SkillRT enables models to avoid these error paths entirely, eliminating redundant interactions and saving substantial tokens.
For weaker models, SkillRT may occasionally consume additional tokens. This occurs because SkillRT improves task completion rates, leading to longer agent execution traces compared to the baseline skill, resulting in modest token increases in edge cases.
Figure 12. Score gain vs. token savings for SkillRT optimization over the Original baseline. Each point is a (model, harness) pair. Color = model tier, shape = harness. Most points land in the upper-right quadrant: better quality and fewer tokens.

### 06 Apr 2026 v2

# SkVM: Compiling Skills for Efficient Execution Everywhere

@@ Abstract.
To address this challenge, we analyze  118,000 skills and draw inspiration from traditional compiler design. We treat skills as code and LLMs as heterogeneous processors. To make portability actionable, we decompose a skill’s requirements into a set of primitive capabilities, and measure how well each model-harness pair supports them. Based on these capability profiles, we propose SkVM, a compilation and runtime system designed for portable and efficient skill execution. At compile time, SkVM performs capability-based compilation, environment binding, and concurrency extraction. At runtime, SkVM applies JIT code solidification and adaptive recompilation for performance optimization.
@@ 1. Introduction
Inspired by how classical computing systems handle code, we introduce SkVM, a compilation and runtime system for skills that enables efficient cross-model execution and provides unified underlying support for skill execution. Through SkVM, skills are compiled into forms tailored to different models, allowing each model to better understand and execute skills. Furthermore, SkVM provides a unified runtime environment that systematically manages skill loading, parsing, and concurrent execution.
@@
We structure SkVM around classical compilation techniques: interpreted execution (jansen2007interpretation, ), ahead-of-time (AOT) compilation (serrano2021javascript, ; serrano2018javascript, ), and just-in-time (JIT) optimization (aho2006compilers, ; aycock2003jit, ; suganuma2000overview, ; ishizaki2000study, ; cramer1997compiling, ). Currently, agents handle skills using only interpreted execution, feeding raw skill text directly to the model. This approach hampers skill portability and execution efficiency. In contrast, SkVM further applies AOT and JIT compilation to optimize for diverse backend models and execution environments:
@@
AOT compilation: Skills inherently encode requirements of model capabilities, environment configuration, and parallelism opportunities. Therefore, SkVM analyzes these requirements before execution and optimizes skills accordingly. Capability-based compilation characterizes the capability gap between LLMs and skill requirements by extracting 26 primitive capabilities. Each capability abstracts a distinct dimension of model behavior with multiple proficiency levels. The compiler measures the target model against these capability dimensions and adapts the skill specification to better align with model strengths and limitations. Second, environment binding extracts implicit dependencies on packages and tools from skill descriptions and generates setup scripts executable at load time, ensuring all dependencies are satisfied before skill execution. Finally, concurrency extraction draws inspiration from classical compiler optimizations for data-level parallelism (DLP) (kalathingal2016dynamic, ), instruction-level parallelism (ILP) (ranganathan1998empirical, ; kastner2001ilp, ), and thread-level parallelism (TLP) (redstone2003mini, ). It extracts parallelism opportunities at similar three granularities from skills and explicitly exposes them to the agent harness, thereby improving task execution efficiency.
@@
JIT optimization: At runtime, SkVM leverages runtime information to continuously optimize skill performance and correctness. Code solidification: Skills may contain parameterized script templates which may be instantiated multiple times during different executions. JIT compilation materializes these high-frequency templates into instantiated executable code, bypassing LLM parsing. Adaptive recompilation: JIT monitoring tracks model execution and adaptively recompiles skills when capability gaps emerge, ensuring skills can self-improve through iterative optimization.
@@ 3. Design Overview
SkVM is a compilation and runtime system for skills. It makes skills portable across heterogeneous targets by combining an AOT compiler (aho2006compilers, ) that specializes skills at install time with a runtime that resolves execution-time uncertainty. Figure 5 shows the architecture.
Figure 5. SkVM architecture. The AOT compiler produces optimized skill variants at install time through three passes. The runtime manages variant selection, JIT optimization, and resource-aware scheduling during execution.
@@ 4. Skill Compilation
When a user installs a skill for the first time, SkVM compiles it against the target (model, harness, host environment) before the skill ever runs. The compiler reads the raw skill text, identifies the target, and produces one or more optimized skill variants through three sequential passes. Each pass closes a different kind of gap between what the skill expects and what the target provides.
@@ 4.1. Capability-Based Compilation
The first pass addresses the model and harness mismatch. Because targets vary widely (§2.5), the compiler cannot embed optimization logic specific to each target. Instead, SkVM introduces primitive capabilities, an abstract vocabulary for expressing what a skill requires and what a target can provide. The compiler profiles each target offline with microbenchmarks for these capabilities, compares the profiling result with the skill’s capability requirements, and selects the optimization strategy based on the gap.
@@ 4.1.2. Measuring the Gap Between Skill and Target
On the target side, SkVM profiles the target with microbenchmarks generated for each primitive capability to measure how well the target supports it. Success on a microbenchmark for a primitive capability at a given level indicates that the model has attained at least that level of proficiency for that capability. This profiling is performed when the user sets up the agent and model. The results are cached and reused for all skill compilations for the same target.
@@ 4.3. Concurrency Extraction
The compiler uses LLM-assisted analysis to decompose the skill into discrete steps, identifying for each step what inputs it consumes and what outputs it produces. It then builds a workflow DAG: nodes are steps, and an edge from step A to step B exists when B consumes an output of A. The compiler also analyzes each step internally for sub-operations that can proceed concurrently. Figure 7 illustrates this process on an incident-response skill. For each parallel opportunity, whether between steps or within a step, the compiler selects a concurrency primitive based on the task structure and the harness. SkVM defines three levels of parallelism.
@@ 5. Skill Runtime with JIT Optimization
AOT compilation generates multiple skill variants for different targets. Consequently, at runtime, SkVM loads the skill variant compiled for the current target. Since AOT compilation is performed only once per skill and target (model+harness) pair, the compilation overhead is modest.
However, some problems only surface at execution time, such as capability gaps that static profiling missed or resource contention among parallel sub-agents competing for rate-limited APIs. Other opportunities, like code patterns that recur across invocations, only become visible after repeated runs. The SkVM runtime addresses both through JIT optimization (§5.2) and a resource-aware scheduler (§5.3).
@@ 5.2. JIT Optimization
Beyond AOT compilation, which tailors primitive capabilities to each target, SkVM further applies JIT optimization to enhance execution quality and efficiency through two complementary mechanisms: adaptive recompilation and code solidification.
@@ 5.2.1. Adaptive Recompilation
Figure 9. Skill compilation effectiveness. Cell values represent task completion rates after SkVM optimization, while cell color indicates the score improvement (green) or regression (red) relative to original baseline skills.
@@ 5.2.2. Code Solidification
In the third stage, the template is instantiated into a callable function: the template code is wrapped with parameter extraction and output handling to produce a standalone shell script or code function. This instantiation step is lightweight and mechanical, as the AOT stage has already done the analysis. Subsequent invocations bypass the LLM entirely: the runtime extracts parameters from the task context, calls the instantiated function, and returns the result. If the solidified code execution causes an agent task to fail or encounters a runtime exception, SkVM triggers a fallback mechanism that re-enables LLM-based code generation to ensure correctness.
@@ 5.3. Resource-Aware Parallel Scheduling
SkVM dynamically monitors system state and schedules task execution accordingly, bridging the gap between compile-time opportunity and run-time reality. For API-bound workloads, it watches response latencies and HTTP 429 (rate limit) signals. For compute-bound workloads, it monitors CPU and memory usage. When pressure rises above a threshold, SkVM applies two mechanisms to mitigate resource contention. First, it throttles the launch of new sub-agents to avoid introducing additional concurrency pressure. Second, for sub-agents that are already running, SkVM selectively suspends a subset of them (i.e., pauses their agent-loop execution), based on per-agent resource usage or launch order, to reduce active competition for shared resources. SkVM also records the effective concurrency of each skill from its previous run on the current system and uses it as a concurrency hint for subsequent executions.
@@ 6. Evaluation
Figure 10. Average task score by skill variant. Four variants are compared across eight models and three harnesses: No Skill, Original, Skill-Creator, and SkVM-Optimized. Tier boundaries (SOTA / Mid / Small) are shaded. SkVM-Optimized skills consistently achieve the highest scores, with the largest absolute gains for weaker models and on OpenClaw.
@@ 6.1. Setup
Models. We evaluate SkVM across eight models spanning three capability tiers to comprehensively assess its effectiveness across varying model capacities. SOTA: claude-opus-4.6 (claude4_6_opus, ), deepseek-v3.2 (deepseek2025v3, ). Mid-tier: gemini-3-flash (gemini_flash, ), qwen3.5-397b (qwen3_5_397b, ), qwen3.5-122b (qwen3_5_122b, ), claude-3.5-haiku (claude3_5_haiku, ). Small: qwen3-30b (qwen3_30b, ), devstral-small (devstral_small_2507, ).
Baselines. We compare SkVM against three baselines. No Skill: the agent executes without any skill loaded. Original: the original skill. Skill-Creator: the skill optimized by Anthropic’s Skill-Creator (anthropic-skill-creator, ) using claude-opus-4.6.
Harnesses. We select three different open-source agent harnesses to evaluate SkVM performance across diverse execution environments. Each harness offers distinct functionality: BareAgent: a minimal harness that injects the compiled skill directly into the system prompt without additional abstraction layers. OpenCode (opencode, ): a full-featured code agent harness with batch tool dispatch capabilities and a rich tool set for code execution. OpenClaw (openclaw, ): a general-purpose agent harness with integrated functional modules designed to support complex multi-step task execution.
Methodology. We first profile the primitive capabilities of different models, and then apply SkVM’s AOT and JIT compilation to generate optimized skill variants tailored to each model. To support the compiled skill artifacts, we extend the agent harnesses with custom loading and management mechanisms that enable SkVM’s optimizations. For each task, we generate five diverse input instances and measure the average task completion rate, token consumption, and end-to-end execution latency. All experiments are conducted on a Mac Mini M4 with 16GB memory.
@@ 6.2. Skill Compilation Effectiveness
We first evaluate SkVM’s effectiveness in improving task completion rates. We compare SkVM-optimized skills against unoptimized baseline skills across different tasks, models, and agent harnesses. For SkVM, the primary gains stem from capability-based compilation and JIT optimization. Figure 9 presents compilation effectiveness across eight models and three harnesses. Cell values represent task completion rates after SkVM optimization, while cell color indicates the score improvement (green) or regression (red) relative to unoptimized baseline skills.
Figure 9 demonstrates SkVM’s optimization effectiveness (compared with original skills) across diverse models and agent harnesses. Weaker models benefit more substantially from SkVM optimization. This finding indicates that for most agent tasks, weaker models possess sufficient capability for the logical components but lack proficiency in non-logical aspects, such as generating complex JSON structures or managing environment dependencies. SkVM’s compilation optimization addresses these gaps through primitive capability refinement, thereby improving overall task completion rates. For stronger models, improvements are more modest, as these models already handle tasks competently; SkVM optimization primarily reduces token consumption and execution latency. Figure 10 aggregates average task scores across all four skill variants for each model and harness. SkVM-optimized skills achieve the highest scores on every model–harness combination. The improvement over the Skill-Creator baseline grows with decreasing model capability: on BareAgent, SkVM-optimized skills improve over Skill-Creator by 25% for qwen3-30b and 10% for devstral-small. In terms of regressions, SkVM-compiled skills reduce task completion rates on only 4.5% of tasks, compared with 15% for the original skills. Using the original skill, OpenCode and OpenClaw differ by up to 13 points, depending on the model, which shows that harness behavior alone materially changes outcomes. SkVM optimization raises scores on both harnesses and reduces this cross-harness gap to at most 5 points.
@@ 6.3. Staged Optimization Breakdown
In 11 of 14 categories, the original skill performs worse than using no skill. After SkVM’s AOT compilation, the average task score improves by 88%. JIT optimization identifies additional capability defects that were not exposed during AOT compilation, further improving performance. After one JIT round, 8 of 14 skills achieve full scores on the evaluation tasks, and after three rounds, 10 of 14 do so.
@@ 6.4. Token and Cost Efficiency with Compilation
Beyond improving task success rates, SkVM provides substantial cost benefits for SOTA LLMs by reducing token consumption. LLMs can correct execution errors through iterative agent loops, refining actions based on environment feedback. Therefore, even when tasks eventually succeed, this approach may incur substantial token waste. Figure 12 shows how task success rates and token costs vary across different model tiers and harnesses after SkVM compilation.
Our evaluation reveals that for most models and harness combinations, SkVM compilation simultaneously improves task success and reduces token consumption. For the strongest model and weakest harness pairing (DS-v3.2 + BareAgent), we observe token savings approaching 40%. This improvement arises because agent loops incur interaction overhead when correcting environment-related and tool-invocation errors; multiple retries are often required. By employing JIT compilation, SkVM enables models to avoid these error paths entirely, eliminating redundant interactions and saving substantial tokens.
For weaker models, SkVM may occasionally consume additional tokens. This occurs because SkVM improves task completion rates, leading to longer agent execution traces compared to the baseline skill, resulting in modest token increases in edge cases.
Figure 12. Score gain vs. token savings for SkVM optimization over the Original baseline. Each point is a (model, harness) pair. Color = model tier, shape = harness. Most points land in the upper-right quadrant: better quality and fewer tokens.